OpenStudy (anonymous):

x2+2x-8=0 solve each equation by the zero factor property

OpenStudy (anonymous):

Factor the equation.

OpenStudy (anonymous):

\[(x+4)(x-2)=0\] Then FOIL (First, Outside, Inside, Last) the equation to check that it is correct. \[(x+4)(x-2)=(x^2+2x-8)\] Then set each factor equal to zero \[(x+4)=0\]\[(x-2)=0\] Solve from there